Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel

P69-002: Optical polarization observations of IXPE targets

PI: Pouya M. Kouch

Multiwavelength polarization, and in particular X-ray polarization, is an important probe of the magnetic field geometry, acceleration physics, and high-energy emission processes of blazar jets. The Imaging X-ray Polarimetry Explorer (IXPE) was successfully launched on the 9th of December 2021 opening a new window to the Universe through high-energy polarization. Here we propose simultaneous multi-band polarization observations of IXPE blazar targets. The proposed observations will supplement our planned multiwavelength campaign in radio, X-rays, and gamma-rays. Comparison of the millimeter-wave, optical, and X-ray polarization and measurement of the mm–optical–UV–X-ray continuum spectra of the blazars, will provide data needed to determine, for the first time, the mechanism and location of the X-ray emission, as well as test particle acceleration models in blazar jets.


ToO triggers

The program aims to provide contemporaneous optical polarization coverage of IXPE observations. We require triggered multi-band (BVRI) polarimetric observations within a day of the IXPE observations. Time-lags of 1-2 days between the optical and X-ray observations are possible if weather or other telescope constrains prevent quasi-simultaneous observations. Longer lags will, however, hamper the achievement of our scientific objectives. For 4 high variable IXPE sources, we require 2 additional observations within two hours to evaluate potential intra-night variability at the time of the IXPE observations. For the intra-night variability we require only VRI observations. The IXPE observing schedule will be known 2-4 weeks in advance.
Calibrations: Twilight flats + alfosc-calibs

The OB generator will be used to request all ToO observations. Scripts to perform all the requested operations will be automatically created and placed in the directory ~/scripts/69-002/. An email will be sent automatically to the observer, which contains all the information about the requested observations, the names and locations of the scripts, instructions for observing and comments from the creator of the trigger.
To summarise:

Go to the directory
  cd ~/scripts/69-002
List the available scripts
  ls -lrt

and execute the OB generated scripts requested in the trigger email.


Date UT window Duration [hh:mm] Comments
06-04-202400:30-[SAT] 00:10ixpe_mrk501
07-04-202400:30-[SAT] 00:10ixpe_mrk501
08-04-202400:30-[SAT] 00:10ixpe_mrk501
09-04-202400:30-[SAT] 00:10ixpe_mrk501
11-04-2024[EAT]-23:00 00:12ixpe_rgb0710
12-04-2024[EAT]-23:00 00:12ixpe_rgb0710
13-04-2024[EAT]-23:00 00:12ixpe_rgb0710
14-04-2024[EAT]-23:00 00:12ixpe_rgb0710
26-04-2024[EAT]-01:0000:10ixpe_mrk421_20240426
27-04-2024[EAT]-01:0000:10ixpe_mrk421_20240427
28-04-2024[EAT]-01:0000:10ixpe_mrk421_20240428
29-04-2024[EAT]-01:0000:10ixpe_mrk421_20240429, Heintz
30-04-2024[EAT]-01:0000:10ixpe_mrk421_20240430
02-05-2024[EAT]-01:0000:10ixpe_mrk421_20240502
03-05-2024[EAT]-01:0000:10ixpe_mrk421_20240503
04-05-2024[EAT]-01:0000:10ixpe_mrk421_20240504
05-05-2024[EAT]-01:0000:10ixpe_mrk421_20240505
06-05-2024[EAT]-01:0000:10ixpe_mrk421_20240506, Stockholm Obs school
07-05-2024[EAT]-01:0000:10ixpe_mrk421_20240507, Stockholm Obs school
08-05-2024[EAT]-01:0000:10ixpe_mrk421_20240508, Stockholm Obs school
09-05-2024[EAT]-01:0000:10ixpe_mrk421_20240509
10-05-2024[EAT]-01:0000:10ixpe_mrk421_20240510, Fynbo
11-05-2024[EAT]-01:0000:10ixpe_mrk421_20240511, Fynbo
12-05-2024[EAT]-01:0000:10ixpe_mrk421_20240512, Fynbo
13-05-2024[EAT]-01:0000:10ixpe_mrk421_20240513
14-05-2024[EAT]-01:0000:10ixpe_mrk421_20240514
17-05-2024[EAT]-01:0000:10ixpe_mrk421_20240517
26-05-2024[EAT]-03:0000:13ixpe_1es1426_20240526
27-05-2024[EAT]-03:0000:13ixpe_1es1426_20240527
28-05-2024[EAT]-03:0000:13ixpe_1es1426_20240528
29-05-2024[EAT]-03:0000:13ixpe_1es1426_20240529, Jones
30-05-2024[EAT]-03:0000:13ixpe_1es1426_20240530, Jones
25-06-2024 [EAT]-01:30 00:13 ixpe_1es1426_20240625
08-07-2024 [EAT]-00:30 00:13 ixpe_1es1426_20240708
08-07-2024 [EAT]-03:00 00:10 ixpe_mrk501_20240708
09-07-2024 [EAT]-03:00 00:10 ixpe_mrk501_20240709
10-07-2024 [EAT]-03:00 00:10 ixpe_mrk501_20240710

Program Tracking
Proposal 69-002
Semester P69: 20240401 - 20241001
Principal Investigator Pouya M. Kouch
Programme Multi-band polarization observations of IXPE targets
Instrument ALFOSC
Time Allocation 30 obs. total, 22 min/obs.
Total Nights 11.0h
Type Mon
Notes Fixed dates to be scheduled ~2 weeks in advance
Total Time Spent 03:47 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
69-002 2024-04-06 00:07 Mon Mrk501 ALHd060224 - ALHd060228 The target was observed successfully. The pointing was correct in the first acquisition. Observed
69-002 2024-04-07 00:09 Mon Mrk501 ALHd070122 - ALHd070127 Observed in clear skies with seeing around 1.0" Observed
69-002 2024-04-08 00:07 Mon Mrk501 ALHd080148 - ALHd080152 Observed in clear skies with seeing around 2.0" Observed
69-002 2024-04-09 00:08 Mon Mrk501 ALHd090235 - ALHd090239 Target observed successfully, although the seeing was very poor. Observed
69-002 2024-04-11 00:09 Mon RGB_0710+591 ALHd110143 - ALHd110147 RGB_0710+591 successfully observed in dusty conditions but within constraints. Observed
69-002 2024-04-12 00:09 Mon RGB_0710+591 ALHd120266 - ALHd120270 Calima, but clear skies. The seeing was at around 1.5 arcsec Observed
69-002 2024-04-13 00:08 Mon RGB_0710+591 ALHd130001 - ALHd130005 Observed during astronomical twilight as the first target of the night. The air was dusty and wind speed was just below the closing limit (12 m/s). Nevertheless, the observation was successful, though a slight offset was present. The object is the topmost one in the frames. Observed
69-002 2024-04-14 00:08 Mon RGB_0710+591 ALHd140017 - ALHd140021 The target was observed successfully on the second attempt. In the first attempt, I tried to optimize the location of the FOV, but I got lost. I considered it a safe option to just observe the target with the same offset as yesterday because I knew the target would be inside the frame, although at the top. Seeing was around 1.8''-2.0''. Observed
69-002 2024-04-26 00:08 Mon Mkn_421_offset, Mrk421 ALHd260273 - ALHd260277 Observed in clear skies with seeing around 0.8" Observed
69-002 2024-04-27 00:00 Mon Could not open dome due to bad weather Failed
69-002 2024-04-28 00:00 Mon Failed because telescope could not be opened due to high humidity and subzero temperatures. Failed
69-002 2024-04-29 00:05 Mon Mkn_421_offset, Mrk421 ALHd290155 - ALHd290159 Observed
69-002 2024-04-30 00:07 Mon Mkn_421_offset, Mrk421 ALHd300060 - ALHd300064 Observed with clear sky and seeing around 1 arcsec. Observed
69-002 2024-05-02 00:08 Mon Mrk421, Mkn_421_offset ALHe020182 - ALHe020186 Observed in clear skies with seeing around 0.9" Observed
69-002 2024-05-03 00:07 Mon Mrk421, Mkn_421_offset ALHe030051 - ALHe030055 Observed in clear skies with seeing below 1.0" Observed
69-002 2024-05-04 00:08 Mon Mrk421, Mkn_421_offset ALHe040054 - ALHe040058 Observed in thin clouds with seeing around 0.8" Observed
69-002 2024-05-05 00:00 Mon Mrk421 ALHe050080 - ALHe050087 Thin clouds, with some holes of clear skies. Seeing of around 0.6 arcsec Observed
69-002 2024-05-06 00:00 Mon -- -- Target not observed due to technical issues with ALFOSC. Failed
69-002 2024-05-07 00:10 Mon Mrk421, Mkn_421_offset ALHe070059-ALHe070063 Observed with very thin clouds and seeing ~0.7". Observed
69-002 2024-05-08 00:10 Mon Mrk421, Mkn_421_offset ALHe080058-ALHe080062 Observed with thin clouds and seeing ~1.00" Observed
69-002 2024-05-09 00:08 Mon Mkn_421_offset, Mrk421 ALHe090163, ALHe090226 - ALHe090229 All targets observed with clear skies or thin clouds. Seeing around 1 arcsec. Observed
69-002 2024-05-10 00:10 Mon Mrk 421 ALHe100146 - ALHe100149 - ALHe100150 Observed with clear skies and subarcsec seeing Observed
69-002 2024-05-11 00:13 Mon Mrk 421 ALHe110099 - ALHe110102 - ALHe110103 there was an error about the ALFOSC focus in the first attempt of the photometry OB, but the second attempt worked Observed
69-002 2024-05-12 00:08 Mon Mrk 421 ALHe120264 - ALHe120267, ALHe120268 - ALHe120268 observed without technical problems Observed
69-002 2024-05-13 00:07 Mon Mrk421, Mkn_421_offset ALHe130052 - ALHe130056 Target observed with clear skies, 1.5" seeing Observed
69-002 2024-05-14 00:09 Mon Mrk421, Mkn_421_offset ALHe140062 - ALHe140066 Observed with clear skies, roughly 1" seeing. Observed
69-002 2024-05-17 00:09 Mon Mrk421, Mkn_421_offset ALHe170248 - ALHe170252 Observed with thin clouds and seeing around 1 arcsec. Observed
69-002 2024-05-26 00:00 Mon ALFOSC was down due to technical problems. Failed
69-002 2024-05-27 00:08 Mon 1ES_1426+428 ALHe270094 - ALHe270098 Observed in clear skies with seeing below 1.0" Observed
69-002 2024-05-28 00:08 Mon 1ES_1426+428 ALHe280251 - ALHe280255 Observed in clear skies with seeing around 1.0" Observed
69-002 2024-05-29 00:09 Mon 1ES_1426+428 ALHe290116 - ALHe290120 Observed with clear skies and sub-arcsecond seeing Observed
69-002 2024-05-30 00:10 Mon 1ES_1426+428 ALHe300087 - ALHe300091 Observed with clear skies and seeing 1-1.5arcsec Observed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: June 11 2024